OpenStudy (micheletheresam):

An airplane needs to reach a forward velocity of 203 km/h to take off. On a 2000 m runway, what is the minimum uniform acceleration necessary for the plane to take flight if it starts from rest?

OpenStudy (micheletheresam):

@gerryliyana

OpenStudy (micheletheresam):

You here? @gerryliyana

OpenStudy (anonymous):

Should the answer be written in m/s2 or km/h2 ?

OpenStudy (micheletheresam):

m/s2

OpenStudy (anonymous):

ok

OpenStudy (micheletheresam):

lmk

OpenStudy (anonymous):

use vt^2 = vo^2 + 2 a. s

OpenStudy (anonymous):

vt^2 = vo^2 + 2 a. s 56.3889^2 = 0^2 + 2 . a . 2000 find a

OpenStudy (anonymous):

are you still here @micheletheresam ?

OpenStudy (anonymous):

vt^2 = vo^2 + 2 a. s 56.3889^2 = 0^2 + 2 . a . 2000 3179.7 = 4000a a = 3179.7/4000 a = 0.795 m/s2

OpenStudy (anonymous):

got it?

OpenStudy (micheletheresam):

Yes thank you very much

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

203 km/h = 56.3889 m/s

OpenStudy (anonymous):

you are awesome @gerryliyana

OpenStudy (anonymous):

@JamesWolf: do you agree with @gerryliyana?